The cost, in millions of dollars, to move x% of pollution in a lake modeled by
C= 6,000/200 - 2x.
Question: What is the cost to remove 75% of the pollutant?

Answer by checkley77(12844)

C= 6,000/200 - 2x.
C=6,000/200-2*.75
C=30-1.5
C=28.5 MILLION DOLLARS.